# Early Classification of Time Series: Taxonomy of Methods and Extensive Benchmark

### v2

@@ 3.2 Anticipation-based methods
Finally, Calimera (?) uses anticipation about the future from another perspective. Instead of trying to guess the likely continuation of $\mathbf{x}_{t}$ which allows one to compute expected future costs, and therefore to wait until there seems no better time to make a prediction, their method is based on predicting directly the difference in cost between predicting the class now or wait at least one more time step. If this difference is positive, then it is better to postpone the prediction. They advocate furthermore, that a calibration step should intervene above the regression in order to make a decision.
3.3 Reinforcement Learning based methods
To learn an ECTS function, it is possible to use an agent that learns what to do by exploring the outcomes associated with different decision strategies as it interacts with the world. The ECTS problem can therefore be recast as a reinforcement learning (RL) problem.
In RL, an agent must learn to associate an action $a$ with each observable state $s$ so that the expected gain be maximized. Let us suppose that the agent-environment interactions break naturally into subsequences which we call episodes
At each time step $t$, the agent perceives the environment’s state $s_{t}$ (i.e. $\mathbf{x}_{t}$), choses an action $a_{t}$ (i.e. either make a prediction now and measure the gain $G_{t}=-\mathrm{C}_{m}(\hat{y}|y)-\mathrm{C}_{d}(\hat{t})$, or postpone the decision and receive the next measurement $x_{t+1}$) according to its current policy $\pi$, where $\pi(a|s)=p(a_{t}=a|s_{t}=s)$ is the probability that action $a$ is taken given the state $s$.
The goal is for the agent to learn an optimal policy $\pi^{\star}$ from sequences of interactions with its environment $\langle s_{t},a_{t},s_{t+1}\rangle$. This can be done by computing the utility function $Q(s,a)$ defined for all (state, action) pairs. By definition:
$Q_{\pi}(s,a)\;=\;\operatornamewithlimits{\mathbb{E}_{\>\pi}^{\>}}\left[G_{t}|S_{t}=s,A_{t}=a\right]$ (12)
and the optimal policy can be derived from:
$Q^{\star}(s,a)\;=\;\operatornamewithlimits{max}_{\pi}Q_{\pi}(s,a)$ (13)
It suffices at each observed state $s$ to chose action $a^{\star}$ such as:
$a^{\star}\;=\;\operatorname*{arg\,max}_{a}Q^{\star}(s,a)$ (14)
One way to learn the function $Q^{\star}$ is using the Q-learning algorithm (?, ?) [^5] Note that other approaches are possible in the reinforcement learning scenario, like learning the utility function $V(s)$ and using TD-learning, or even learning the policy $\pi$ directly. The Q-learning approach is however widely used. and its variants for continuous definition of environment’s states, such as Deep Q-learning (?).
In End-to-end approaches, only one function is responsible for both stopping and making a prediction about the class of $\mathbf{x}_{t}$ whereas, in separable approaches involves two combined functions, respectively dedicated to triggering the prediction and to the classification itself. In the case of Reinforcement Learning based approaches, this distinction takes the following form:
  1. 1.
    The separable approach. RL can be used to learn the Trigger function only, once the classifiers $h_{t}$ have been learnt. In that case, the set of actions $\mathcal{A}_{t}$ at each time step $t$ is restricted to two elements: {‘decide now’, ‘postpone decision’}. The $Q$ function evaluates for each state the expected gain for each of the two possibilities, allowing one to decide what to do at each time step.
  2. 2.
    The End-to-End approach. RL can also be used to learn at once both when to trigger a prediction and what prediction to make. In principle, it suffices to extend the set of actions to $\mathcal{A}_{t}=\{`\textit{postpone decision'},c_{1},\ldots,c_{N}\}$, where there are $N$ classes. Either, the agent postpones the decision, or it predicts a class for the incoming time series $\mathbf{x}_{t}$.
In the literature, Reinforcement Learning-based ECTS approaches frequently include Deep Learning modules:
Separable RL approaches:
The Earliest (Early and Adaptive Recurrent Label ESTimator) uses a RNN architecture (?) to make the prediction and a Reinforcement Learning agent trained jointly using policy gradient, to trigger prediction or not. If a prediction is triggered, the hidden representation given by the RNN is sent to a Discriminator, whose role is to predict a class, given this representation. The model has been adapted to deal with irregularly sampled time series (?). ? (?) extend the ECTS framework to channel filtering, using here also Reinforcement Learning.
End-to-end RL approaches:
? (?, ?) use a Deep Q-Network (?), alongside a specifically designed reward signal, encouraging the agent to find a good trade-off between earliness and accuracy. Those types of approaches also naturally extend to online settings where time series are not of fixed length. (?) introduce EarlytStop-RL, in which model-free RL is used to address the problem of early diagnosis of lung cancer.
3.4 Deep Learning based methods
Alongside the Reinforcement Learning based approaches, there exist deep learning methods that do not use RL.
The Decouple ETSC Network (Detscnet) (?) architecture leverages a gradient projection technique in order to jointly learn two sub-modules: one for variable-length series classification, the other for the early exiting task.
The End-to-end Learned Early Classification of Time Series method (Elects) leverages a LSTM architecture, adding a stopping prediction head to the network and adapting the loss function to promote good early predictions (?).
@@ 4 Experiments & Results
This section presents the extensive set of experiments carried out in order to provide a consistent and fair evaluation of a wide range of the existing literature’s methods. We first describe the experimental protocol used. We then turn to the experiments and their results. Figure 4 provides a synthetic view of the organization of these experiments.
@@
  • In Section 4.2, the main state-of-the-art and the three baseline methods are evaluated using a widely used cost setting, i.e. with a a binary balanced misclassification cost and a linear delay cost.
  • In Section 4.3, methods are tested in an anomaly detection scenario where the misclassification cost matrix is severely imbalanced, with false negatives being much more costly than false positives, and where the delay cost is no longer linear with time but increases exponentially with time.
  • Finally, Section 4.4 briefly describes a set of other experimental setups, derived from either standard setting or the anomaly detection one, including for instance testing the impact of z-normalization. Complementary results can be found in Appendix B.
@@
Figure 2: Experiments diagram: each line corresponds to one (or two) full benchmark runs, representing in total twelve full benchmarks. While this Section mainly discusses results about the cost settings in Subsections 4.2 and 4.3, many other alternative experiments are briefly analyzed in Subsection 4.4 and are more detailed in the Appendix B.
The experiments presented here aim to evaluate the effects of design choices on method performance and thus to provide answers to the questions:
  • Do anticipation-based methods perform better than blind or myopic ones?
  • Do methods that are cost-informed for their decision (i.e. explicitly estimating costs) perform better than methods that are cost-uninformed (e.g. confidence-based) (see Section 2.3.2).
  • How the various methods fare when modifying the form of the delay cost and/or the misclassification cost matrix?
@@ 4.1 Experimental Protocol
This section covers the shared part of the experimental protocol for all experiments irrespective of the choice of the cost functions (see Sections 4.2 and 4.3 for this).
@@ 4.1.1 Evaluation of the performance
The ECTS problem explicitly balances the pressure to make a correct prediction and a time pressure. The correctness of the prediction is measured by the misclassification cost $\mathrm{C}_{m}(\hat{y}|y)$ where $\hat{y}$ is the prediction and $y$ is the true class. The time pressure is sanctioned by a delay cost $\mathrm{C}_{d}(t)$ that is assumed to be positive and, in most applications, an increasing function of time. Sometimes, the two can be combined when the misclassification cost is function of the time : $\mathrm{C}_{md}(\hat{y}|y,t)$.
Therefore, for each test times series $\mathbf{x}^{i}$, an ECTS method incurs a cost assumed to be of the following additive form: $\mathrm{C}_{m}(\hat{y}_{i}|y_{i})\,+\,\mathrm{C}_{d}(\hat{t})$, where $\hat{t}$ is the time when the system decided to make a prediction, this prediction being $\hat{y}_{i}$.
For a test set of $M$ time series, the average cost of a method is:
@@
This is accordingly the criterion with which we evaluated the methods in the experiments reported in this paper.
In addition, in order to assess how the methods adapt to various balances between the misclassification and the delay costs, we vary the settings of these costs by weighting them during training and testing. The performance of the methods is therefore evaluated using the weighted average cost, as defined in Equation 16, for different values of the costs balance $\alpha$, ranging from 00 to $1$, with a $0.1$ step:
$AvgCost_{\alpha}=\frac{1}{N}\sum_{i=0}^{N}\alpha\times C_{m}(\hat{y}_{i}|y_{i})+(1-\alpha)\times C_{d}(\hat{t}_{i})$ (16)
@@
4.1.2 Optimization of the parameters of the methods
@@
4.1.3 Comparing the trigger methods
In order to carry out a fair comparison between the tested methods, we isolated as far as possible the triggering function, responsible for deciding when to stop receiving measurements, from the prediction one, responsible for predicting a label to the incoming time series. As advocated by ? (?), we have chosen the MiniROCKET algorithm (?) to be the base classifier for all methods. It is indeed recognized as among the best performing classifier in the Time Series classification literature as well as one of the fastest one. Of course, distinguishing the decision component from the prediction one is only possible for the “separable” methods. In our experiments, we chose to not evaluate “end-to-end” methods, leaving that for future work.
Trigger models: Nine trigger models were selected from the literature based on their usage and their performances [^6] The EDSC algorithm (?), even though available in the provided library, is not included in the following experiments, due to high space and time complexity (which hinders fair comparisons).
  • Economy-$\gamma$-Max (?): triggers a decision if the predicted cost expectation is the lowest at time $t$ when compared with the expected cost for all future time steps (cf. Section 3.2, Anticipation-based).
  • Calimera (?): triggers a decision when a regressor model which predicts the difference between the current observed cost and the minimum cost in the future is negative (cf. Section 3.2, Anticipation-based).
  • Stopping Rule (?): uses a trigger function based on a linear combination of confidence estimates and a delay measure linear on time (cf. Section 3.1, Confidence-based).
[…]
  • Ecec (?): defines a confidence measure, based on the aggregated confidence of the predictions up to time $t$, and triggers a prediction if it exceeds a threshold, tuned by explicit grid-search (cf. Section 3.1, Confidence-based).
  • Ecdire (?): determines “safe” timestamps, based on classifier performance, from which predictions about possible classes can be made. Predictions cannot be triggered if those timestamps have not been reached. In addition, the difference between the two highest predicted probabilities must also exceed a certain threshold. (cf. Section 3.1, Confidence-based).
  • Ects (?): computes the first time $t$ for which nearest neighbors of the incoming time series $\mathbf{x}_{t}$ in the training set were given a label that did not change by the classifier (cf. Section 3.2, Anticipation-based).
All these methods have been re-implemented using Python, reproducing results close to the published ones. Except the code for the Ects implementation, which has been taken from ? (?). Hyper-parameters are the ones chosen in the original published methods. Code to reproduce the experiments is available publicly at https://github.com/ML-EDM/papers_experiments/tree/main/ects_reloading.
@@ 4.1.3 Comparing the trigger methods
  • Alap (As Late As Possible) always waits the complete series to trigger the prediction.
  • Proba Threshold is a natural, confidence-based, cost-informed at training time, baseline: it triggers a prediction if the estimated probability of the likeliest prediction exceeds some threshold, found by grid search (cf. Section 3.1, Confidence-based).
4.1.4 Calibration of the classifications
Like ? (?), we add a calibration step when learning the classifiers, i.e. Platt’s scaling (?). Indeed, as we are dealing with collections of independently trained classifiers, the prediction scores may not remain consistent with one another over the time dimension. However, the trigger methods usually have their parameters set with the same values for all time steps. This is the case, for example with the Proba Threshold approach. In addition, some approaches such as Calimera and Economy-$\gamma$-Max exploit the estimated posterior probabilities $\{p(y|\textbf{x}_{t})\}_{y\in\mathcal{Y}}$ to estimate the future cost expectation. It is therefore highly desirable for all classifiers, at all times, to have their output calibrated.
@@ 4.1.5 Datasets and training protocol
Datasets [^7] All original datasets of the paper can be downloaded, already prepared and splitted, from https://urlz.fr/qRqu: In order to be able to directly compare our results to past experiments, we first use the usual TSC datasets from the UCR Archive (?) with the default split. In total, we have used 77 datasets from the UCR Archive, i.e. the ones with enough training samples to satisfy our experimental protocol end-to-end (blue cylinder in Figure 4). In this way, most of the datasets used by either ? (?) and ? (?) or by ? (?) are contained in our experiments.
A second collection of non z-normalized data sets is also provided. In this way, the associated potential information leakage is avoided (see Section 1). Any difference with the performance obtained on the z-normalized data sets can thus signal the danger of z-normalization with firm evidence. Considering the limited amount of non z-normalized datasets within the UCR archive (?), we have decided to look for complementary new datasets so as to provide another collection of datasets. To this end, the Monash archive for extrinsic regression (?), provided 20 new time series datasets, for which we have discretized the the numerical target variable into binary classes based on a threshold value. For instance, if this threshold is equal to the median value of the regression target, the resulting classification datasets will be balanced in term of classes (as in Section 4.2). Note that this threshold can be chosen differently to get imbalanced datasets (as in Section 4.3.2), several thresholds could also be used to increase number of classes. As a result, we get a new set of classification tasks, as has recently been done by ? (?). Finally, 35 datasets have been gathered: 15 from the original archive and 20 from the Monash extrinsic regression archive. (orange cylinder in Figure 4).
@@
When not using predefined splits, the train sets are split into two distinct sets in a stratified fashion: a first one to train the different classifiers, corresponding to 60% of the training set and another one to train the trigger model, trained over the 40% left. The set used to train the classifiers, is itself split into two different sets in order to train calibrators, using 30% of the given data. Because of this procedure, we have been led to exclude some of the datasets, due to their limited training set size.
All the experiments have been performed using a linux operating machine, with an Intel Xeon E5-2650 2.20GHz (24-cores) and 252GB of RAM. Proceeding all datasets (including both blue and orange cylinders) over all competing approaches takes between 9-10 days, using MiniROCKET classifier, which is the most efficient tested.
@@ 4.2.1 Cost definition
The misclassification cost is symmetrical and balanced, the delay cost is linear. They can be defined as follows:
@@
Thus, for each dataset, the AvgCost is bounded between 0 and 1.
@@ 4.2.2 Results and analysis
For comparability reasons, this first set of experiments is analyzed over the classical ECTS benchmark used in the literature so far (blue cylinder in Figure 4). Results over the new, non z-normalized, datasets can be found in Appendix B.
@@
Figure 3(b) provides a global views about the relative performances of the tested methods. The Wilcoxon-Holm Ranked test provides an overall statistical analysis. It examines the critical difference among all techniques to plot the method’s average rank in a horizontal bar. Lower ranks denote better performance, and the methods connected by a horizontal bar are similar in terms of statistical significance. When evaluated by their average rank on all data sets with respect to the average cost (Equation 16), here for $\alpha=0.5$, four methods significantly outperform the others:
        Methods Confidence Anticipation Cost informed
Stopping Rule train
Proba Threshold train
Economy-$\gamma$-max train & test
Calimera train
Figure 3(a) allows a closer look, this time varying the relative costs of misclassification and delaying prediction using Equation 16, where a small value of $\alpha$ means that delay cost is paramount. $90\%$ level confidence intervals have been computed using bootstrap [^8] Resample with replacement has been done a large number of times (10.000 $\times$) and are reported as shaded colors on the figure. The statistic of interest is studied, here the mean, by examining the bootstrap distribution at the desired confidence level.. Again, the same four methods top the others for almost every values of $\alpha$. Not surprisingly, the baseline Asap (predict as soon as possible) is very good when the delay cost is very high, while Alap (predict at time $T$) is very good when there is no cost associated with delaying decision.
It is remarkable that, in this cost setting, the simple Proba Threshold method exhibits a strong performance for almost all values of $\alpha$. It is therefore worth including in the evaluation of new methods. However, while Figures 3(a), 3(b) are useful for general analysis, they do not provide insights about how the Accuracy vs Earliness trade-off is optimized for each of the competitor. Figure 4 provides some explanation for this.
Figure 4: Pareto front, displaying for each $\alpha$ the Accuracy on the $y$ axis and $Earliness$ on the $x$ axis. Best approaches are located on the top left corner. In zoomed boxes, on the right of the Figure, points corresponding to a single $\alpha$ are highlighted, while other points are smaller and gray. Each of the trigger model is optimizing the trade-off in its own way, resulting in many different approaches having points in the Pareto dominant set.
In this figure, the two evaluation measures: ‘Accuracy’ and ‘Earliness’, are considered as dimensions in conflict. The Pareto front is the set of points for which no other point dominates with respect to both Accuracy and Earliness. It is drawn here when varying their relative importance using $\alpha$ (in the set $\{0,0.1,0.2,\ldots,1.0\}$).
@@
Each of the leading methods Stopping Rule, Proba Threshold, Economy and Calimera have at least one point on the Pareto front and generally exhibits a combined performance very close to it. A closer look reveals how each approach optimizes the earliness vs. accuracy trade-off differently for a fixed cost. If we consider $\alpha=0.8$, for example, it appears that Economy takes its decision earlier than Proba Threshold, itself being more precocious than Ecec. Because this is also an area of problems where the delay cost is low, by doing so, Economy prevents itself from benefiting from waiting for more measurements and increasing its performance. Hence its slight downward slope on Figure 3(a) for high values of $\alpha$.
It is worth noting that the two naive baselines Asap and Alap perform better than the majority of approaches on seven $\alpha$ values out of ten. This is specially the case when the delay cost is large, i.e. for $\alpha\in[0.1,0.3]$, for which the Asap baseline is as competitive as top performers. Globally, the performance of Proba Threshold is remarkable in this cost setting. Even though, it is simply based on a single threshold on the confidence in the current prediction, its performance makes it one of the best methods.
The results computed over the proposed datasets ensemble (i.e. orange cylinder) are displayed in Figure 8 of Appendix B. No significant changes can be observed in the ranking of competing approaches.
One question is how such simple method, like Proba Threshold, can adapt to scenarios where the misclassification and delay costs, not being symmetrical for the misclassification cost, and not linear for the delay one, reflect other application settings.

### v6

@@ 3.2 Anticipation-based methods
$\bullet$ Calimera (Bilski & Jastrzebska, 2023) uses anticipation about the future from another perspective. Instead of trying to guess the likely continuation of $\mathbf{x}_{t}$ which allows one to compute expected future costs, and therefore to wait until there seems no better time to make a prediction, their method is based on predicting directly the difference in cost between predicting the class now and the best reachable cost from next timestep to last one $T$. If this difference is positive, then it is better to postpone the prediction. They advocate furthermore, that a calibration step should intervene on the collection of classifiers, in order to build meaningful regression target to learn the trigger model.
@@ 4 Experiments & Results
This section presents the extensive set of experiments carried out in order to provide a consistent and fair evaluation of a wide range of existing literature’s methods. We first describe the experimental protocol used. We then turn to the experiments and their results. Figure 2 provides a synthetic view of the organization of these experiments.
@@
  • In Section 4.2, eight state-of-the-art methods and the three baseline ones are evaluated using a widely used cost setting, i.e. with a binary balanced misclassification cost and a linear delay cost.
  • In Section 4.3, methods are tested in an anomaly detection scenario [^4] We consider anomalies to be actual phenomena of interest, such as the failure of a machine (Boniol et al., 2024). These anomalies may be accompanied by revealing precursor signals in the time series, which an ECTS system should be able to detect by optimizing both the accuracy of prediction and its earliness. where the misclassification cost matrix is severely imbalanced, with false negatives being much more costly than false positives, and where the delay cost is no longer linear with time but increases exponentially with time.
  • Finally, Section 4.4 briefly describes a set of other experimental setups, derived from either the standard setting or the anomaly detection one, including, for instance, testing the impact of z-normalization. Complementary results can be found in Appendix LABEL:app_results.
@@
Figure 2: Experiments diagram. While this section mainly discusses results about the cost settings in Subsections 4.2 and 4.3, many other alternative experiments are briefly analyzed in Subsection 4.4 and are more detailed in the Appendix LABEL:app_results. Details on the used datasets can be found in Appendix E.
The experiments presented here aim to evaluate the effects of design choices on method performance and thus provide answers to the questions:
  • Do anticipation-based methods perform better than myopic ones?
  • Do methods that are cost-informed for their decisions (i.e. explicitly estimating costs) perform better than methods that are cost-uninformed? (see Section 2.2.1)
  • How do the various methods fare when modifying the form of the delay cost and/or the misclassification cost matrix?
@@ 4.1 Experimental Protocol
This section covers the shared part of the experimental protocol for all the experiments, irrespective of the choice of the cost functions (see Sections 4.2 and 4.3 for this). Additional implementation specifications can be found in Appendix B.
@@ 4.1.1 Evaluation of the performance
[…]
For a test set of $M$ time series, the average cost of a method (Equation 8) is used as the criterion with which to evaluate the methods.
@@
In addition, in order to assess how the methods adapt to various balances between the misclassification and the delay costs, we vary the settings of these costs by weighting them during training and testing. The performance of the methods is therefore evaluated using the weighted average cost, as defined in Equation 9, for different values of the costs balance $\alpha$, ranging from 0 to $1$, with a $0.1$ step:
$AvgCost_{\alpha}=\frac{1}{M}\sum_{i=0}^{M}\alpha\times C_{m}(\hat{y}_{i}|y_{i})+(1-\alpha)\times C_{d}(\hat{t}_{i})$ (9)
@@
4.1.2 Comparing the trigger methods
Our experiments aim foremost at comparing the trigger methods used that are responsible for deciding when to make a prediction about the class of the incoming time series. To this end, all compared methods in our experiments use the same prediction component. As advocated by Bilski & Jastrzebska (2023), we have chosen the MiniROCKET algorithm (Dempster et al., 2021) to be the base classifier for all methods. It is indeed recognized as among the best performing classifiers in the time series classification literature as well as one of the fastest ones. However, we have carried out additional experiments with two other classifiers (see Section 4.4). In our experiments, we have additionally reported results for end-to-end methods in order to situate the performance of these methods in respect to separable ones in Appendix C.2.
@@ 4.1.2 Comparing the trigger methods
All these methods have been re-implemented using Python, reproducing results close to the published ones. Except for the code for the Ects implementation, which has been taken from Kladis et al. (2021). Hyperparameters are the ones chosen in the original published methods. Code to reproduce the experiments is available publicly at (see https://github.com/ML-EDM/ECTS_survey) .
@@
  • Alap (As Late As Possible) always waits for the complete series to trigger the prediction.
  • Proba Threshold is a natural, confidence-based, cost-informed, baseline: it triggers a prediction if the estimated probability of the likeliest prediction exceeds some threshold, found by grid search (cf. Section 3.1, Confidence-based).
@@ 4.1.3 Datasets and training protocol
Two collections of datasets [^6] All original datasets of the paper can be downloaded, already prepared and split, from https://urlz.fr/qRqu: In order to be able to directly compare our results to past experiments, we first use the usual TSC datasets from the UCR Archive (Dau et al., 2019) with the default split. In total, we have used 77 datasets from the UCR Archive, i.e. the ones with enough training samples to satisfy our experimental protocol from the start to the end, i.e. with at least one example per class within each of the used disjoint subsets (see Splitting strategy below). (blue cylinder in Figure 2). In this way, most of the datasets used by either Mori et al. (2017a) and Lv et al. (2019), or by Achenchabe et al. (2021) are contained in our experiments.
A second collection of non z-normalized datasets is also provided. In this way, the associated potential information leakage is avoided (see Section 1). Any difference in the performance obtained on the z-normalized data sets can thus signal the danger of z-normalization with firm evidence. Considering the limited amount of non z-normalized datasets within the UCR archive (Dau et al., 2019), we have decided to look for complementary new datasets so as to provide another collection of datasets. To this end, the Monash archive for extrinsic regression (Tan et al., 2020), provided 20 new time series datasets, for which we have discretized the numeric target variable into binary classes based on a threshold value. For instance, if this threshold is equal to the median value of the regression target, the resulting classification datasets will be balanced in terms of classes (as in Section 4.2). Note that this threshold can be chosen differently to get imbalanced datasets (as in Section 4.3.2), several thresholds could also be used to increase the number of classes. As a result, we obtain a new set of classification tasks, as has recently been done by Middlehurst et al. (2024b). In the end, 34 datasets have been gathered: 14 from the original archive and 20 from the Monash extrinsic regression archive. (orange cylinder in Figure 2). More details about the selection of datasets can be found in Appendix B.4
@@
When not using predefined splits, the train sets are split into two distinct sets in a stratified fashion: a first one to train the different classifiers, corresponding to 40% of the training set and another one to train the trigger model, trained over the 60% left. The set used to train the classifiers is itself split into two different sets in order to train calibrators, using 30% of the given data. Because of this procedure, we have been led to exclude some of the datasets, due to their limited training set size.
All the experiments have been performed using a Linux operating system, with an Intel Xeon E5-2650 2.20GHz (24-cores) and 252GB of RAM. Proceeding all datasets (including both blue and orange cylinders) over all competing approaches takes between 9-10 days, using MiniROCKET classifier, which is the most efficient tested.
@@ 4.2.1 Cost definition
The misclassification cost is symmetrical and balanced; the delay cost is linear. They can be defined as follows:
@@
Thus, for each dataset, the $\textit{AvgCost}_{\alpha}$ is bounded between 0 and 1, as, within this cost definition, the average misclassification (resp. temporal) cost, across examples is equivalent to the 1Accuracy1-Accuracy (resp. Earliness) measure, with $Accuracy=\frac{1}{M}\sum_{i=1}^{M}\mathbb{1}(\hat{y}_{i}=y_{i})$ and $Earliness=\frac{1}{M\times T}\sum_{i=1}^{M}\hat{t}_{i}$.
@@ 4.2.2 Results and analysis
For comparability reasons, this first set of experiments is analyzed over the classical ECTS benchmark used in the literature so far (blue cylinder in Figure 2). Results over the new, non z-normalized, datasets can be found in Appendix LABEL:app_results.
@@
When evaluated by their average rank on all data sets with respect to the average cost (Equation 9), here for $\alpha=0.5$, four methods significantly outperform the others:
Table 2: Leading ECTS methods and their properties along dimensions underlined in the taxonomy.
Methods Confidence Anticipation Cost-informed
Stopping Rule
Proba Threshold
Economy-$\gamma$-max ✓(test)
Calimera
Figure 3(b) provides a view about the relative performances of the tested methods in terms of the average cost induced using the methods for $\alpha=0.5$. The Wilcoxon-Holm Ranked test provides an overall statistical analysis. It examines the critical difference among all techniques to plot the method’s average rank in a horizontal bar. Lower ranks denote better performance, and the methods connected by a horizontal bar are similar in terms of statistical significance.
It is remarkable that, in this cost setting, the simple Proba Threshold method exhibits a strong performance for almost all values of $\alpha$. It is therefore worth including in the evaluation of new methods. However, while Figures 3(a), 3(b) are useful for general analysis, they do not provide insights about how the Accuracy vs. Earliness trade-off is optimized for each of the competitors. Figure 4 provides some explanation for this.
Figure 4: Pareto front, displaying for each $\alpha$ the Accuracy on the $y$-axis and EarlinessEarliness on the $x$-axis. Best approaches are located on the top left corner. In zoomed boxes, on the right of the Figure, points corresponding to a single $\alpha$ are highlighted, while other points are smaller and gray. Each of the trigger model is optimizing the trade-off in its own way, resulting in many different approaches having points in the Pareto dominant set.
In this figure, the two evaluation measures: Accuracy and Earliness, are considered as dimensions in conflict. The Pareto front is the set of points for which no other point dominates with respect to both Accuracy and Earliness. It is drawn here when varying their relative importance using $\alpha$ (in the set $\{0,0.1,0.2,\ldots,1.0\}$).
[…]
Each of the leading methods Stopping Rule, Proba Threshold, Economy and Calimera have at least one point on the Pareto front and generally exhibit combined performance very close to it. A closer look reveals how each approach optimizes the Earliness vs. Accuracy trade-off differently for a fixed cost. If we consider $\alpha=0.8$, for example, it appears that Economy takes its decision earlier than Proba Threshold, itself being more precocious than Ecec. Because this is also an area of problems where the delay cost is low, by doing so, Economy prevents itself from benefiting from waiting for more measurements and increasing its performance. Hence its slight downward slope on Figure 3(a) for high values of $\alpha$.
It is worth noting that the two naive baselines Asap and Alap perform better than the majority of approaches on seven $\alpha$ values out of ten. This is especially the case when the delay cost is large, i.e. for $\alpha\in[0.1,0.3]$, for which the Asap baseline is as competitive as top performers. Globally, the performance of Proba Threshold is remarkable in this cost setting. Even though it is simply based on a single threshold on the confidence in the current prediction, its performance makes it one of the best methods.
The results computed over the proposed datasets ensemble (i.e. orange cylinder) are displayed in Figure LABEL:fig:std_orange_cylinder of Appendix LABEL:app_results. No significant changes can be observed in the ranking of competing approaches.